OpenStudy (anonymous):

Let f(x)=arctan(squareroot(4x−1)). Find f′(x)

OpenStudy (ash2326):

We have \[f(x)=\tan ^{-1} (\sqrt {4x-1})\] We know \[\frac{d}{dx} \tan^{-1} x= \frac{1}{1+x^2}\] and also we know \[\frac{d}{dx} f(g(x)= f'(g(x)) \times g'(x)\] so Let's begin \[f'(x)= \frac{1}{1+ (\sqrt{4x-1})^2} \frac{d}{dx} (\sqrt{4x-1})\] \[f'(x)= \frac{1}{1+ (\sqrt{4x-1})^2} \times \frac{1}{2} \frac{1}{\sqrt{4x-1}}\times \frac{d}{dx}(4x)\] \[f'(x)= \frac{1}{1+ (\sqrt{4x-1})^2} \times \frac{1}{2} \frac{1}{\sqrt{4x-1}}\times 4\] We get now \[f'(x)= \frac{1}{1+ {4x-1}} \times \frac{1}{2} \frac{1}{\sqrt{4x-1}}\times 4\] \[f'(x)= \frac{1}{{4x}} \times \frac{1}{2} \frac{1}{\sqrt{4x-1}}\times 4\] \[f'(x)= \frac{1}{{\cancel 4x}} \times \frac{1}{2} \frac{1}{\sqrt{4x-1}}\times \cancel4\] We get finally \[f'(x)= \frac{1}{2x} \frac{1}{\sqrt{4x-1}}\]